OpenStudy (anonymous):

Complete each arithmetic sentence to show the distributive property. a.8(6+4) b.10(a+b)

OpenStudy (jiteshmeghwal9):

8(6+4)=8(10)=>80

OpenStudy (jiteshmeghwal9):

now use distributive property to proof if we gt 80 as the answer

OpenStudy (jiteshmeghwal9):

8(6)+8(4) 48+32 80

OpenStudy (jiteshmeghwal9):

yes ! we r getting the same number hence we can solve this type of question using distributive property.

OpenStudy (jiteshmeghwal9):

can u prove the next arithmetic sentence @de4354 ???

OpenStudy (anonymous):

ok so 10(a+10)b 10x10=100 right?????

OpenStudy (jiteshmeghwal9):

nope

OpenStudy (jiteshmeghwal9):

\[\LARGE{\color{red}{Distributive \space property :}}\]\[\LARGE{\color{green}{a(c+b)=a(c)+a(b) \implies ac+ab}}\]

OpenStudy (jiteshmeghwal9):

so, 10(a+10)b 10(a)+10(10)b 10a+100b

OpenStudy (anonymous):

SO 10(a+b)= 10(a)+10(B)

OpenStudy (jiteshmeghwal9):

yes!

OpenStudy (anonymous):

OK thanks i truly didn't get that

OpenStudy (jiteshmeghwal9):

so now u have gt it :)